In a large introductory statistics lecture​ hall, the professor reports that 60​% of the students enrolled have never taken a calculus​ course, 20​% have taken only one semester of​ calculus, and the rest have taken two or more semesters of calculus. The professor randomly assigns students to groups of three to work on a project for the course. You are assigned to be part of a group. What is the probability that of your other two​ groupmates, ​a) neither has studied​ calculus? ​b) both have studied at least one semester of​ calculus? ​c) at least one has had more than one semester of​ calculus?
Answer #1

a) Probability that neither of the two of our group members have studied calculus is computed here as:

= 0.62 = 0.36

Therefore 0.36 is the required probability here.

b) Probability that both have studied at least one semester of calculus

= ( 1 - 0.6)2 = 0.42 = 0.16

Therefore 0.16 is the required probability here.

c) Probability that at least one has had more than one semester of calculus is computed here as:

= 1 - Probability that both have less than or equal to 1 semester of calculus

= 1 - (0.6 + 0.2)2 = 1 - 0.64 = 0.36

Therefore 0.36 is the required probability here.